40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Sep 20 03:12:14 2011, Anamaria, Update, SUS, Jenne's Scripts started 
    Reply  Tue Sep 20 04:12:26 2011, Jenne, Update, SUS, Jenne's Scripts started 
       Reply  Tue Sep 20 09:44:44 2011, Jenne, Update, SUS, Jenne's Scripts started 
          Reply  Tue Sep 20 14:53:13 2011, Jenne, Update, SUS, Jenne's Scripts started 
Message ID: 5477     Entry time: Tue Sep 20 09:44:44 2011     In reply to: 5476     Reply to this: 5479
Author: Jenne 
Type: Update 
Category: SUS 
Subject: Jenne's Scripts started 

Quote:

Quote:

I followed Jenne's instructions, ran the matrix filler script and then set the optics to freeswing. Someone has to burt resture and damp them in the morning.

 Thanks!  I'll give them a little more time, then restore things.

 I began restoring the optics at ~9:30am, so I have a full 6 hours of data, in case I need that much to separate the Pos/Side modes on some of the optics.  They are all damping again with their original matricies.

ELOG V3.1.3-